Cable exerts a constant upward tension of magnitude 1.98 x 10^4 N on a 1.80 x 10^4 kg elevator as it rises through a vertical distance of 4.10 m

1. What is the work done by the tension force on the elevator.
2. What is the work done by the force of gravity on the elevator.
1
Expert's answer
2020-03-23T10:38:24-0400

1. The work done by the tension force on the elevator is given by


"W_1=Fd=1.98\times 10^4\times 4.10=8.12\times 10^4\:\rm J."

2. The work done by the force of gravity on the elevator is given by


"W_2=mgd=1.80\times 10^4\times 9.8\times 4.10=7.23\times 10^5\:\rm J."